OpenStudy (anonymous):

an automobile accelerates from rest at 1+ 3* sqrt (t) mph/sec for 9 seconds. What is the velocity after 9 seconds?

OpenStudy (anonymous):

1+3sqrt(9) 1+3(3) 1+9 10mph

OpenStudy (xishem):

^ Wouldn't this find the acceleration at time t, not the velocity?

OpenStudy (xishem):

*time 9s

OpenStudy (anonymous):

Good point Xishem... I wondered about that too, but I can't tell from the question X(

OpenStudy (xishem):

Would integrating the function from 0 to 9 give the correct answer?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Hm, when the problem states it like that... does it mean that the velocity is the function 1 + 3*(t) or the acceleration?

OpenStudy (anonymous):

the function is of acceleration

OpenStudy (anonymous):

so its integral will be velocity

OpenStudy (anonymous):

It means acceleration....yeah, Xishem, you're right, this will require integration

OpenStudy (anonymous):

Yeah, the book says mph/sec.. how should it be stated if it was giving the velocity?

OpenStudy (anonymous):

(t+6t^(3/2)) from 0-->9

OpenStudy (anonymous):

i would change either the hours to seconds or vice versa

OpenStudy (anonymous):

so your answer can be in the form mps or mph

OpenStudy (anonymous):

this can be done by dimensional analysis

OpenStudy (dumbcow):

i get v(t) = t + 2t*sqrt(t) v(9) = 9 +2*9*sqrt(9) = 63 mph

OpenStudy (anonymous):

integral of sqrt (t) = [2t^(3/2)]/3 Multiply that by the 3 constant.. = 2t^(3/2) am I doing something wrong? I can't get sqrt (t)

OpenStudy (anonymous):

it's the same thing... 2t*t^(1/2)=2t^(3/2)

OpenStudy (anonymous):

Oh I see. I would have never though of to do that... makes calculations so much easier, haha.